OpenStudy (konradzuse):

evaluate each integral (x^2 - 1) / (x) dx

OpenStudy (anonymous):

x-1/x x^2/2-lnx+c

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac {x^{2}-1}{x} dx = \int\limits_{}^{} x dx - \int\limits{}^{}\frac{1}{x}dx\]

OpenStudy (anonymous):

=1/2 x^2 -ln(x) + C